A hydrocarbon contains 10.5g of carbon per gram of hydrogen. 1 litre of the vapour of the hydrocarbon at 127oC and 1 atmosphere pressure weighs 2.8g. find the molecular formula.

Grade:10

1 Answers

Jitender Pal
9 years ago
Given P = 1 atmV = 1L, T = 127oC = 127 + 273 = 4000K
PV = nRT (Ideal gas equation)
Or n =PV/RT = 1 *1/0.082 *400 = 0.0304
Mol. wt Mass/Moles = 2.8/0.0304 = 92.10
Element
Wt. of element
Relation no. of atom
Ratio of of atoms
Whole no. of atoms
C
H
10.5
1.0
10.5/12
= 0.875
1.0/1 = 1
0.875/0.875
= 1
1/0.875 = 1.14
1 * 7 = 7
1.14 *7 = 8
∴ Emperical formula = C7H
Emperical formula, wt = 12 * 7 + 1 * 8 = 92
n = Molecular wt/Emperical formula wt = 92.10/92 = 1
molecular formula = n * empirical formula
= 1 (C7H8) = C7H8